# Tac-Man: Tactile-Informed Prior-Free Manipulation of Articulated Objects

### v3

@@ IV-C Manipulation under Imperfect Priors
Imperfect priors present a significant challenge in robotic manipulation, even when the articulation structure is known. Errors in model parameters can make these priors unreliable, as demonstrated by the example of an off-axis door, depicted in Fig. 7(a). Unlike conventional doors with hinges located along one edge, the off-axis door features a hinge positioned away from the edge, complicating the acquisition of accurate priors.
The impact of imperfect priors is illustrated through a challenging manipulation task, in which the robot is programmed to follow an arc trajectory to open a microwave oven with a revolute joint, as shown in Fig. 7(b). However, a $10~{}\%$ error in the trajectory’s radius introduces a small but critical imperfection in the priors. This error, without any real-time corrective adjustments, accumulates a loss that eventually leads to the gripper losing contact with the handle, resulting in task failure.
In contrast, Fig. 7(c) demonstrates Tac-Man’s approach, employing adaptive adjustments during the manipulation to effectively counteract the inaccuracies in the priors. By dynamically compensating for the unforeseen deviation, Tac-Man ensures alignment with the correct interaction direction. This adaptability allows Tac-Man to successfully complete the task, fully opening the microwave oven door, showcasing its efficacy in overcoming the limitations imposed by imperfect priors.
@@ IV-D Manipulation under Unknown Priors
Encountering objects whose articulation mechanisms defy accurate perception introduces a scenario where priors on kinematics remain unknown. Such conditions, characterized by complex movements that cannot be succinctly represented by conventional kinematic models like prismatic or revolute joints, present a significant challenge. Existing research has scarcely addressed the manipulation of objects requiring combined translational and rotational motions due to the difficulty in modeling these actions explicitly. For example, Fig. 8(a) illustrates an object designed compactly, necessitating that its handle executes both linear and rotational movement simultaneously. Another example is shown in Fig. 8(b), where a vise handle follows a spiral path. The ability to accurately identify and model such complex motion patterns remains beyond the reach of current perception methods.
@@ IV-E Manipulation under Obsolescent Priors
Dynamic environments often introduce unpredictable perturbations, such as human interventions or environmental turbulence, rendering initially accurate priors for object kinematics obsolescent. To evaluate Tac-Man’s performance in this context, we conducted experiments depicted in Fig. 9 in which the position of the articulated object was deliberately altered during manipulation. This was achieved by applying force to one side of the object, leading to a shift in either the rotation axis of a cabinet (Fig. 9(a)) or the opening direction of a drawer (Fig. 9(b)), as indicated by the hand icons.
@@ V Large-scale Verification in Simulation
Given the logistical challenges of physically experimenting with a comprehensive variety of objects, we look to simulation studies to further validate the generalization capability of Tac-Man. Utilizing NVIDIA Isaac Sim, our simulations cover a wide range of objects, mirroring the setup of our real-world experiments as outlined in Sec. V-A. For a thorough assessment, we incorporate objects with prismatic and revolute joints from the PartNet-Mobility dataset [36], reflecting the common articulations found in existing datasets.
@@
To showcase the effectiveness of Tac-Man, we select several representative prior-based and prior-free methods as baselines, as detailed in Sec. V-B. In Sec. V-C, we present both qualitative and quantitative results from our simulations. The findings compellingly demonstrate Tac-Man’s proficiency in manipulating objects with a wide range of articulations, highlighting its adaptability and superior generalization across diverse scenarios compared to the baseline methods.
@@ V-A Environmental Setup
Our simulation environment, depicted in Fig. 10, features a robot arm equipped with a gripper poised to engage an object’s handle. Initially, the robot is programmed to follow a preliminary direction $\boldsymbol{R}^{1}$, applying Tac-Man to articulate the object. A critical aspect of Tac-Man is its reliance on the compliance of silica gel material to accurately track contact between the gripper and the handle. To mimic this in a rigid body simulation environment, we employ specialized designs for contact measurement, detailed in Sec. -A.
The simulation divides test objects into two categories: standard objects with prismatic or revolute joints featuring various handle designs from existing datasets, and randomly created custom objects with complex trajectories. From the PartNet-Mobility dataset [36], we select objects with detailed articulation annotations provided by GAPartNet [38], as outlined in Sec. -B. For the evaluation, $25$ objects with prismatic joints and another $25$ with revolute joints were pre-processed. A manipulation is deemed successful if a revolute joint rotates beyond $60.0^{\circ}$or a prismatic joint extends over $250.0~{}\mathrm{mm}$. Fig. 10(a) showcases a subset of these objects and their initial grasps.
To assess Tac-Man’s adaptability beyond basic joint types, we also evaluate it on articulated objects designed with randomly generated trajectories. These are constructed by sampling Bézier curve segments of various orders ($2$, $3$, $4$, and $5$) on a $2$D plane, ensuring no self-intersections. Each curve inspires a $3$D playboard that defines a manipulation path, with a toy train at the start point acting as the handle. Physical collisions are configured to restrict the handle’s movement along the trajectory, with the construction process detailed in Sec. -C. Fig. 10(b) illustrates an example of a $3$-order Bézier curve, its corresponding play-board, and the simulation’s initial state. Additional examples of generated play-boards are presented in Fig. 10(c), highlighting the increased complexity and manipulation challenge with higher-order curves. For the evaluation, $25$ curves of orders $2$ to $5$ are sampled, each tested once. Success is achieved when the gripper moves the toy train from one end of the curve to the other.
@@ V-B Selected Baselines
FlowBot3D, introduced by Eisner et al. [19], is a recent advancement using modern learning techniques for generalized manipulation based on visual feedback. It employs a flow prediction model to predict part mobility and compute actions accordingly. It should be noticed that, to align with our simulation settings, we apply only the flow prediction model and action computation strategy. We assess its performance using the same tasks described in Sec. V-A as those used for Tac-Man, without additional fine-tuning.
@@ V-C Results
Our simulation results substantiate the efficacy of the setup in evaluating Tac-Man. In Fig. 11, we highlight five representative examples: (a–c) illustrate the manipulation of objects with standard prismatic and revolute joints, while (d–e) detail scenarios involving intricate trajectory manipulations. The direction of pull and subsequent adjustments are indicated by the orange arrows. The simulated tactile feedback is visualized in the grey area beneath each figure, in which inactive, reference and current contact points are depicted as grey dots, blue circles, and red dots, respectively. Orange arrows, enlarged five-fold for clarity, depict the displacement from reference to current markers, with their size increasing in proportion to the error magnitude.
@@
Consistent with our real-world experiments, deviations in the manipulation direction lead to deviation (Fig. 11(a-c)) in the tactile pattern, signaling the need for corrective movements beyond the preliminary direction. Once the contact deviation surpasses $\alpha\epsilon$ (Fig. 11(b–e)), Tac-Man adeptly adjusts to re-establish stable contact.
@@
The quantitative results for the three baseline methods and Tac-Man are summarized in Fig. 12. To provide a more informed measure of performance beyond binary success or failure, we introduce a distance-weighted success rate $SR_{W}$, defined as:
@@
where $d_{E}$ represents the expected contact point moving distance required for success, as specified in Sec. V-A, and $d_{A}$ is the actual handle moving distance calculated from the final handle position in each trial.
Fig. 12 presents both mean and standard deviation of the weighted success rate across objects under different categories. Specifically, columns $1$-$2$ display weighted success rates for manipulating objects with prismatic and revolute joints, while columns $3$-$6$ show performance on intricate manipulation trajectories, with PB-$n$ indicating performance on Bézier trajectories of order $n$. Notably, Tac-Man achieves a $100\%$ weighted success rate across all objects with various handle designs, markedly outperforming the three baseline methods and verifying its effectiveness. Key observations for the baseline methods include:
@@ VI-A Advancing Robotic Manipulation with Tac-Man
Our study has validated the efficacy of Tac-Man, a tactile-informed, prior-free approach for manipulating articulated objects. The core of Tac-Man is leveraging tactile feedback to discern discrepancies between the preliminary direction and actual interaction directions, informing dynamic adjustments to ensure precise manipulations. A standout feature of Tac-Man is its operational independence from pre-existing knowledge about object kinematics, presenting a considerable advantage over traditional methods reliant on priors. This capability is particularly beneficial in environments where priors may be ambiguous (Sec. IV-B), imperfect (Sec. IV-C), unknown (Sec. IV-D), or obsolescent (Sec. IV-E)—conditions prevalent in dynamic, human-centric settings. While classical control methods like impedance control can operate without object priors, they may fail with complex trajectories, as shown in Fig. 12.
Tac-Man enhances robotic autonomy by obviating the need for manually inputting accurate object kinematic models, a common requirement in earlier strategies [4, 5, 6]. It sidesteps the limitations of methods that depend on visual perception for determining object kinematics [10, 11, 12, 14, 15, 16, 21], which are vulnerable to the ambiguities and inaccuracies intrinsic to visual data [39], as demonstrated by the reduced effectiveness of pre-planned methods under noise in Fig. 12. Unlike approaches that utilize multi-frame observations to clarify ambiguities [8, 13, 20, 19, 17, 18, 7, 9], Tac-Man does not presuppose the existence of specific articulation types, thus demonstrating its versatility in managing novel articulation mechanisms as evidenced in Sec. IV-D and Fig. 12.
@@
Additionally, Tac-Man showcases remarkable adaptability to dynamic changes within environments, as explored in Sec. IV-E. Its ability to adjust to unforeseen human interventions underscores a vital strength for robots operating in human-centric environments, where static assumptions about object kinematics can lead to manipulation failures. This adaptability is essential for ensuring seamless human-robot coexistence, enabling robots to effectively navigate the inherent unpredictability of real-world settings.
@@ VI-B Limitations and Future Work
Tac-Man’s effectiveness primarily stems from its ability to compensate for contact deviation, which is modeled in a geometric perspective so that both direction and magnitude of deviation can be obtained, as detailed in Sec. III-E. This paper demonstrates the implementation of Tac-Man with GelSight-based sensor, as one economic representative for those sensors with the ability to measure contact geometry. We would like to clarify that the high-resolution property of GelSight is not a major consideration in this work, as indicated in the setup for experiments. Actually, all these sensors can be effectively adopted within the Tac-Man framework and should have a very similar formulation, with only minor variation in the pre-processing procedure where contact deviation geometry is extracted from raw sensor data. For example, the TacTip family [63] uses an array of pins to detect surface deformations, which can be mapped to contact points in a manner similar to our current implementation. Therefore, one important future work will be to investigate whether the extraction of contact deviation geometry with different sensors can be generalized to some degree to improve the succinctness of Tac-Man formulation.
@@
The effectiveness of Tac-Man relies on obtaining a stable initial contact $\mathcal{C}_{1}$ and having a preliminary (rough) direction to manipulate an articulated object. In this study, we manually establish the initial (rough) grasp and preliminary direction for all experiments. Crucially, Tac-Man’s success stems from its ability to adjust control input on the fly based on tactile feedback; hence, this initial grasp need not be precise and can sometimes even be intentionally wrong. To demonstrate Tac-Man’s resilience, we conducted experiments on a drawer (see Fig. 13), providing the robot arm with intentionally incorrect grasps and directions. Despite these inaccuracies, Tac-Man adapts and successfully completes the task of opening the drawer.
@@
The initial grasping force critically influences Tac-Man’s efficiency by directly impacting the stable contact state $\mathcal{C}_{1}$ and associated tactile feedback. While it minimally affects the computation of the optimal transformation matrix in Eq. 23 (which relies on relative changes to $\mathcal{C}_{1}$ for contact stabilization), it significantly influences the calculation of $t$, the maximum distance during the execution stage. This $t$ value, derived from the optimization program (7), depends on absolute changes. Simulation results in Fig. 14 demonstrate that both excessively large and small grasping forces prolong manipulation times. This aligns with intuition: larger forces may exceed the gripper’s maximum elastic deformation capacity, while smaller forces increase slippage risk. Both scenarios are more likely to violate the stable contact constraint (2), reducing the maximum $t$ value from program (7) and necessitating additional iterations between execution and recovery stages. Crucially, within the elastic deformation region, Tac-Man yields successful manipulations of articulated objects despite varied execution time, demonstrating its robustness. Based on these findings, we practically select a grasping force that yields a deformation of approximately half the maximum normal elastic deformation $e_{n}$. While this approach is effective, future work should explore methods for adaptively selecting the optimal grasping force to further enhance performance across various manipulation scenarios.
@@
While the Tac-Man demonstrates notable capabilities in the prior-free manipulation of articulated objects, an area for potential enhancement is its time efficiency. As indicated in Tab. II and Fig. 15, the completion time for manipulation tasks using Tac-Man tends to exceed that of human performance. This disparity is largely attributed to the current implementation of Tac-Man, which operates without the benefits of closed-loop control.
@@
In its present form, Tac-Man employs a cautious, step-by-step exploratory approach. This method, devoid of pre-established object priors or learning, requires careful evaluation of, and response to, tactile feedback. The task duration figures reported in Tab. II and Fig. 15 primarily reflect this initial phase of exploration, critical for the robot to accurately interpret and interact with the object.
@@
Looking ahead, enhancing Tac-Man to include velocity-based control elements could significantly improve its time efficiency. Such a development would enable quicker and more responsive adaptations during initial exploration, potentially reducing the overall time required for manipulations. It would lead to a more balanced approach, achieving manipulation robustness and efficiency on a par with human performance, especially in dynamic environments as discussed in Sec. IV-E.
@@
Tac-Man, as discussed in Sec. II-E, is a model-based tactile servoing method. This approach offers benefits such as reduced data collection requirements and enhanced explainability [84]. However, integrating Tac-Man into a learnable framework represents a promising avenue for future research. This integration aims to generalize the strategy of utilizing tactile sensing in articulated object manipulation to a broader range of challenges, including dexterous manipulation [3], cable following [74], surface tracking [53], and others.

### v4

@@ IV-C Manipulation under Imperfect Priors
Imperfect priors present a significant challenge in robotic manipulation, even when the articulation structure is known. Errors in model parameters can make these priors unreliable, as demonstrated by the example of an off-axis door, depicted in Fig.˜7(a). Unlike conventional doors with hinges located along one edge, the off-axis door features a hinge positioned away from the edge, complicating the acquisition of accurate priors.
The impact of imperfect priors is illustrated through a challenging manipulation task, in which the robot is programmed to follow an arc trajectory to open a microwave oven with a revolute joint, as shown in Fig.˜7(b). However, a $10~\%$ error in the trajectory’s radius introduces a small but critical imperfection in the priors. This error, without any real-time corrective adjustments, accumulates a loss that eventually leads to the gripper losing contact with the handle, resulting in task failure.
In contrast, Fig.˜7(c) demonstrates Tac-Man’s approach, employing adaptive adjustments during the manipulation to effectively counteract the inaccuracies in the priors. By dynamically compensating for the unforeseen deviation, Tac-Man ensures alignment with the correct interaction direction. This adaptability allows Tac-Man to successfully complete the task, fully opening the microwave oven door, showcasing its efficacy in overcoming the limitations imposed by imperfect priors.
@@ IV-D Manipulation under Unknown Priors
Encountering objects whose articulation mechanisms defy accurate perception introduces a scenario where priors on kinematics remain unknown. Such conditions, characterized by complex movements that cannot be succinctly represented by conventional kinematic models like prismatic or revolute joints, present a significant challenge. Existing research has scarcely addressed the manipulation of objects requiring combined translational and rotational motions due to the difficulty in modeling these actions explicitly. For example, Fig.˜8(a) illustrates an object designed compactly, necessitating that its handle executes both linear and rotational movement simultaneously. Another example is shown in Fig.˜8(b), where a vise handle follows a spiral path. The ability to accurately identify and model such complex motion patterns remains beyond the reach of current perception methods.
@@ IV-E Manipulation under Obsolescent Priors
Dynamic environments often introduce unpredictable perturbations, such as human interventions or environmental turbulence, rendering initially accurate priors for object kinematics obsolescent. To evaluate Tac-Man’s performance in this context, we conducted experiments depicted in Fig.˜9 in which the position of the articulated object was deliberately altered during manipulation. This was achieved by applying force to one side of the object, leading to a shift in either the rotation axis of a cabinet (Fig.˜9(a)) or the opening direction of a drawer (Fig.˜9(b)), as indicated by the hand icons.
@@ V Large-scale Verification in Simulation
Given the logistical challenges of physically experimenting with a comprehensive variety of objects, we look to simulation studies to further validate the generalization capability of Tac-Man. Utilizing NVIDIA Isaac Sim, our simulations cover a wide range of objects, mirroring the setup of our real-world experiments as outlined in Sec.˜V-A. For a thorough assessment, we incorporate objects with prismatic and revolute joints from the PartNet-Mobility dataset [36], reflecting the common articulations found in existing datasets.
@@
To showcase the effectiveness of Tac-Man, we select several representative prior-based and prior-free methods as baselines, as detailed in Sec.˜V-B. In Sec.˜V-C, we present both qualitative and quantitative results from our simulations. The findings compellingly demonstrate Tac-Man’s proficiency in manipulating objects with a wide range of articulations, highlighting its adaptability and superior generalization across diverse scenarios compared to the baseline methods.
@@ V-A Environmental Setup
Our simulation environment, depicted in Fig.˜10, features a robot arm equipped with a gripper poised to engage an object’s handle. Initially, the robot is programmed to follow a preliminary direction $\boldsymbol{R}^{1}$, applying Tac-Man to articulate the object. A critical aspect of Tac-Man is its reliance on the compliance of silica gel material to accurately track contact between the gripper and the handle. To mimic this in a rigid body simulation environment, we employ specialized designs for contact measurement, detailed in Sec.˜-A.
The simulation divides test objects into two categories: standard objects with prismatic or revolute joints featuring various handle designs from existing datasets, and randomly created custom objects with complex trajectories. From the PartNet-Mobility dataset [36], we select objects with detailed articulation annotations provided by GAPartNet [38], as outlined in Sec.˜-B. For the evaluation, $25$ objects with prismatic joints and another $25$ with revolute joints were pre-processed. A manipulation is deemed successful if a revolute joint rotates beyond $60.0^{\circ}$or a prismatic joint extends over $250.0~\mathrm{mm}$. Fig.˜10(a) showcases a subset of these objects and their initial grasps.
To assess Tac-Man’s adaptability beyond basic joint types, we also evaluate it on articulated objects designed with randomly generated trajectories. These are constructed by sampling Bézier curve segments of various orders ($2$, 33, 44, and $5$) on a 22D plane, ensuring no self-intersections. Each curve inspires a 33D playboard that defines a manipulation path, with a toy train at the start point acting as the handle. Physical collisions are configured to restrict the handle’s movement along the trajectory, with the construction process detailed in Sec.˜-C. Fig.˜10(b) illustrates an example of a $3$-order Bézier curve, its corresponding play-board, and the simulation’s initial state. Additional examples of generated play-boards are presented in Fig.˜10(c), highlighting the increased complexity and manipulation challenge with higher-order curves. For the evaluation, $25$ curves of orders $2$ to $5$ are sampled, each tested once. Success is achieved when the gripper moves the toy train from one end of the curve to the other.
@@ V-B Selected Baselines
FlowBot3D, introduced by Eisner et al[19], is a recent advancement using modern learning techniques for generalized manipulation based on visual feedback. It employs a flow prediction model to predict part mobility and compute actions accordingly. It should be noticed that, to align with our simulation settings, we apply only the flow prediction model and action computation strategy. We assess its performance using the same tasks described in Sec.˜V-A as those used for Tac-Man, without additional fine-tuning.
@@ V-C Results
Our simulation results substantiate the efficacy of the setup in evaluating Tac-Man. In Fig.˜11, we highlight five representative examples: (a–c) illustrate the manipulation of objects with standard prismatic and revolute joints, while (d–e) detail scenarios involving intricate trajectory manipulations. The direction of pull and subsequent adjustments are indicated by the orange arrows. The simulated tactile feedback is visualized in the grey area beneath each figure, in which inactive, reference and current contact points are depicted as grey dots, blue circles, and red dots, respectively. Orange arrows, enlarged five-fold for clarity, depict the displacement from reference to current markers, with their size increasing in proportion to the error magnitude.
@@
Consistent with our real-world experiments, deviations in the manipulation direction lead to deviation (Fig.˜11(a-c)) in the tactile pattern, signaling the need for corrective movements beyond the preliminary direction. Once the contact deviation surpasses $\alpha\epsilon$ (Fig.˜11(b–e)), Tac-Man adeptly adjusts to re-establish stable contact.
@@
The quantitative results for the three baseline methods and Tac-Man are summarized in Fig.˜12. To provide a more informed measure of performance beyond binary success or failure, we introduce a distance-weighted success rate $SR_{W}$, defined as:
@@
where $d_{E}$ represents the expected contact point moving distance required for success, as specified in Sec.˜V-A, and $d_{A}$ is the actual handle moving distance calculated from the final handle position in each trial.
Fig.˜12 presents both mean and standard deviation of the weighted success rate across objects under different categories. Specifically, columns 11-22 display weighted success rates for manipulating objects with prismatic and revolute joints, while columns 33-66 show performance on intricate manipulation trajectories, with PB-$n$ indicating performance on Bézier trajectories of order $n$. Notably, Tac-Man achieves a $100\%$ weighted success rate across all objects with various handle designs, markedly outperforming the three baseline methods and verifying its effectiveness. Key observations for the baseline methods include:
@@ VI-A Advancing Robotic Manipulation with Tac-Man
Our study has validated the efficacy of Tac-Man, a tactile-informed, prior-free approach for manipulating articulated objects. The core of Tac-Man is leveraging tactile feedback to discern discrepancies between the preliminary direction and actual interaction directions, informing dynamic adjustments to ensure precise manipulations. A standout feature of Tac-Man is its operational independence from pre-existing knowledge about object kinematics, presenting a considerable advantage over traditional methods reliant on priors. This capability is particularly beneficial in environments where priors may be ambiguous (Sec.˜IV-B), imperfect (Sec.˜IV-C), unknown (Sec.˜IV-D), or obsolescent (Sec.˜IV-E)—conditions prevalent in dynamic, human-centric settings. While classical control methods like impedance control can operate without object priors, they may fail with complex trajectories, as shown in Fig.˜12.
Tac-Man enhances robotic autonomy by obviating the need for manually inputting accurate object kinematic models, a common requirement in earlier strategies [4, 5, 6]. It sidesteps the limitations of methods that depend on visual perception for determining object kinematics [10, 11, 12, 14, 15, 16, 21], which are vulnerable to the ambiguities and inaccuracies intrinsic to visual data [39], as demonstrated by the reduced effectiveness of pre-planned methods under noise in Fig.˜12. Unlike approaches that utilize multi-frame observations to clarify ambiguities [8, 13, 20, 19, 17, 18, 7, 9], Tac-Man does not presuppose the existence of specific articulation types, thus demonstrating its versatility in managing novel articulation mechanisms as evidenced in Sec.˜IV-D and Fig.˜12.
@@
Additionally, Tac-Man showcases remarkable adaptability to dynamic changes within environments, as explored in Sec.˜IV-E. Its ability to adjust to unforeseen human interventions underscores a vital strength for robots operating in human-centric environments, where static assumptions about object kinematics can lead to manipulation failures. This adaptability is essential for ensuring seamless human-robot coexistence, enabling robots to effectively navigate the inherent unpredictability of real-world settings.
@@ VI-B Limitations and Future Work
Tac-Man’s effectiveness primarily stems from its ability to compensate for contact deviation, which is modeled in a geometric perspective so that both direction and magnitude of deviation can be obtained, as detailed in Sec.˜III-E. This paper demonstrates the implementation of Tac-Man with GelSight-based sensor, as one economic representative for those sensors with the ability to measure contact geometry. We would like to clarify that the high-resolution property of GelSight is not a major consideration in this work, as indicated in the setup for experiments. Actually, all these sensors can be effectively adopted within the Tac-Man framework and should have a very similar formulation, with only minor variation in the pre-processing procedure where contact deviation geometry is extracted from raw sensor data. For example, the TacTip family [63] uses an array of pins to detect surface deformations, which can be mapped to contact points in a manner similar to our current implementation. Therefore, one important future work will be to investigate whether the extraction of contact deviation geometry with different sensors can be generalized to some degree to improve the succinctness of Tac-Man formulation.
@@
The effectiveness of Tac-Man relies on obtaining a stable initial contact $\mathcal{C}_{1}$ and having a preliminary (rough) direction to manipulate an articulated object. In this study, we manually establish the initial (rough) grasp and preliminary direction for all experiments. Crucially, Tac-Man’s success stems from its ability to adjust control input on the fly based on tactile feedback; hence, this initial grasp need not be precise and can sometimes even be intentionally wrong. To demonstrate Tac-Man’s resilience, we conducted experiments on a drawer (see Fig.˜13), providing the robot arm with intentionally incorrect grasps and directions. Despite these inaccuracies, Tac-Man adapts and successfully completes the task of opening the drawer.
@@
The initial grasping force critically influences Tac-Man’s efficiency by directly impacting the stable contact state $\mathcal{C}_{1}$ and associated tactile feedback. While it minimally affects the computation of the optimal transformation matrix in Eq.˜23 (which relies on relative changes to $\mathcal{C}_{1}$ for contact stabilization), it significantly influences the calculation of $t$, the maximum distance during the execution stage. This $t$ value, derived from the optimization program (7), depends on absolute changes. Simulation results in Fig.˜14 demonstrate that both excessively large and small grasping forces prolong manipulation times. This aligns with intuition: larger forces may exceed the gripper’s maximum elastic deformation capacity, while smaller forces increase slippage risk. Both scenarios are more likely to violate the stable contact constraint (2), reducing the maximum $t$ value from program (7) and necessitating additional iterations between execution and recovery stages. Crucially, within the elastic deformation region, Tac-Man yields successful manipulations of articulated objects despite varied execution time, demonstrating its robustness. Based on these findings, we practically select a grasping force that yields a deformation of approximately half the maximum normal elastic deformation $e_{n}$. While this approach is effective, future work should explore methods for adaptively selecting the optimal grasping force to further enhance performance across various manipulation scenarios.
@@
While the Tac-Man demonstrates notable capabilities in the prior-free manipulation of articulated objects, an area for potential enhancement is its time efficiency. As indicated in Tab.˜II and Fig.˜15, the completion time for manipulation tasks using Tac-Man tends to exceed that of human performance. This disparity is largely attributed to the current implementation of Tac-Man, which operates without the benefits of closed-loop control.
@@
In its present form, Tac-Man employs a cautious, step-by-step exploratory approach. This method, devoid of pre-established object priors or learning, requires careful evaluation of, and response to, tactile feedback. The task duration figures reported in Tab.˜II and Fig.˜15 primarily reflect this initial phase of exploration, critical for the robot to accurately interpret and interact with the object.
@@
Looking ahead, enhancing Tac-Man to include velocity-based control elements could significantly improve its time efficiency. Such a development would enable quicker and more responsive adaptations during initial exploration, potentially reducing the overall time required for manipulations. It would lead to a more balanced approach, achieving manipulation robustness and efficiency on a par with human performance, especially in dynamic environments as discussed in Sec.˜IV-E.
@@
Tac-Man, as discussed in Sec.˜II-E, is a model-based tactile servoing method. This approach offers benefits such as reduced data collection requirements and enhanced explainability [84]. However, integrating Tac-Man into a learnable framework represents a promising avenue for future research. This integration aims to generalize the strategy of utilizing tactile sensing in articulated object manipulation to a broader range of challenges, including dexterous manipulation [3], cable following [74], surface tracking [53], and others.